xx TestReduce

TestReduce

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters 7 Interfaces

General Implements: testreduce

Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 testreducex3gfs 10                    01  compact 11 0
11 10 phixgf3 01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 testreducex2gfs 10                    01  compact 11 0
11 10 phixgf2 01 description 11 Test grid function
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 testreducex1gfs 10                    01  compact 11 0
11 10 phixgf1 01 description 11 Test grid function
11                    10                    01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 testreducexscalars 10                    01  compact 11 0
11 10 phixscalar 01 description 11 Test scalar
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 testreducex3igfs 10                    01  compact 11 0
11 10 phixigf3 01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 INT
11 testreducex2igfs 10                    01  compact 11 0
11 10 phixigf2 01 description 11 Test grid function
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 INT

11

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 testreducex1igfs 10                    01  compact 11 0
11 10 phixigf1 01 description 11 Test grid function
11                    10                    01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 INT
11 testreducexiscalars 10                    01  compact 11 0
11 10 phixiscalar 01 description 11 Test scalar
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 INT

8 Schedule This section lists all the variables which are assigned storage by thorn CactusTest/TestReduce. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always:  11 x
11 testreducex3gfs  11 x
11 testreducex2gfs  11 x
11 testreducex1gfs  11 x
11 testreducex3igfs 11 x
11 testreducex2igfs 11 x
11 testreducex1igfs 11 x
11 testreducexscalars 11 x
11 testreducexiscalars 11 x

11 x  11 x

11

Scheduled Functions

CCTKxINITIAL x testreducexinitial x initialise variables x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxPOSTINITIAL x testreducecxmaximum x test out maximum reductions from c x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxPOSTINITIAL x testreducefxmaximum x test out maximum reductions from fortran x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function

CCTKxPOSTINITIAL x testreducecxsum x test out sum reductions from c x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxPOSTINITIAL x testreducefxsum x test out sum reductions from fortran x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function